How many ounces of the two alcohol solutions in the illustration must be mixed to obtain 100 ounces of a 12.2% solution?
The illustration has two bottles = a bottle
8% + 15% = 100 oz (12.2%)
Please list how the problem was solved. Thank you!

Answer by mananth(16949)

percent ----------------quantity oz
8 ----------------x
15 ----------------100-x
12.20% ----------------100
8x+15(100-x)=12.2*100
8x+1500 -15x=1220
8x-15x=1220-1500
-7x=-280
/-7
x=40 oz 8%
balance 60 oz 15%